OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is greater than 6 and 1 die shows a 3? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

I think this question was answered a couple days ago: Probability of getting a 3 = 1/6 (1 number out of 6) Numbers needed on 2nd die to get 7 or higher: 4, 5, 6 : 1/2 Total probability: 1/6*1/2*2 = 1/6 (multiplied by 2 because dice can switch places)